Calculate the approximate volume of a 0.62 mol sample of gas at 10.6 °C and a pressure of 1.05 atm.


1
Expert's answer
2021-03-26T05:43:07-0400

PV = nRT

V = nRT/P

n = 0.62 mole

R = 0.0821L /atm/mol/K

T = 10 + 273 = 283K

P = 1.05 atm

V = (0.62mol × 0.0821 L/atm/mol/K × 283K)/1.05 atm

= 13.74839 L